OpenStudy (anonymous):

(EARLY ALGEBRA II) bx+a=dx+c Solve for x. State any restrictions on the variables.

OpenStudy (anonymous):

move dx to the right side and a to the left bx-dx=c-a take the x out as a factor of the left side x(b-d)=c-a divid both side by b-d x=(c-a)/(b-d)

OpenStudy (anonymous):

=> b must not equal d

OpenStudy (anonymous):

If b=d you would have division by zero

OpenStudy (anonymous):

@TheMind @ajdirig Thank you guys so much :) Really cleared things up for my. Literal equations have never really been my forte! I appreciate it! :D